OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6

OpenStudy (anonymous):

it is eaither -3 or 6 i think

OpenStudy (anonymous):

Get rid of one variable first: x 6x + 4y = 12 6x + 9y = -18 Subtract and you'll get: -5y = 30 y = -6 Plug into one of the equation: 3x + 2y = 6 3x + 2(-6) = 6 3x -12 = 6 3x = 18 x = 6 Answer: x = 6, y = -6

OpenStudy (experimentx):

simply solve them!! you know how to solve them right??

OpenStudy (anonymous):

OOPS i was kinda positive it was -6 but i guss i was wrong lol